### v1

# Variable Rate Learned Wavelet Video Coding with Temporal Layer Adaptivity

Abstract
Learned wavelet video coders provide an explainable framework by performing discrete wavelet transforms in temporal, horizontal, and vertical dimensions. With a temporal transform based on motion-compensated temporal filtering (MCTF), spatial and temporal scalability is obtained. In this paper, we introduce variable rate support and a mechanism for quality adaption to different temporal layers for a higher coding efficiency. Moreover, we propose a multi-stage training strategy that allows training with multiple temporal layers. Our experiments demonstrate Bjøntegaard Delta bitrate savings of at least -17% compared to a learned MCTF model without these extensions. Our method also outperforms other learned video coders like DCVC-DC. Training and inference code is available at: https://github.com/FAU-LMS/Learned-pMCTF.
Index Terms:
Motion compensated temporal filtering, wavelet transform, learned video compression
I Introduction
The main paradigm in learned compression is nonlinear transform coding [1] similar to predictive transform coding in traditional compression. However, wavelet transforms have desirable properties for compression because their compromise between spatial and frequency resolution fits the correlation structure of visual data. Methods built on learned wavelet transforms also provide an explainable framework, as their latent space has a defined structure corresponding to a wavelet decomposition. In addition, they support lossless compression in contrast to most learned methods. Learned wavelet coding is therefore specifically suited for tasks such as volumetric medical image coding [2, 3, 4] and biomedical video coding [5].
In this paper, we address limitations of existing learned wavelet video coding schemes [6, 7]: The models use a trainable temporal wavelet transform realized using motion-compensated temporal filtering (MCTF). MCTF supports spatial and temporal resolution scaling, where temporal scalability is achieved by a hierarchical group of pictures (GOP) structure with multiple temporal decomposition levels [8, 9]. Up to now, learned MCTF coders require a separate model for every rate-distortion point [6, 7]. This lack of support for continuous rate adaptation leads to increased training and storage costs. Moreover, existing MCTF models only support training on two frames. This results in a train-test mismatch, as during inference the models are evaluated with multiple temporal layers for GOPs larger than two. In addition, training on a single temporal layer does not allow quality adaptions to each layer, which gives considerable coding gains for traditional video coders in the context of B frame coding. We tackle these issues with the following contributions:
  • We enable smooth rate adjustment with trainable quality scaling parameters.
  • We introduce temporal layer adaptive quality scaling considering the layer structure of MCTF for a higher coding efficiency.
  • We propose a multi-stage training strategy that allows training on two temporal decomposition levels. This way, we double the possible training sequence length and reduce the train-test mismatch for MCTF.
Figure 1: Overview of a learned wavelet image coder [7]. $\bm{x}$ is one channel of an image in the YCbCr color space. The subbands $\bm{y}$ obtained from a learned discrete wavelet transform (DWT) are coded sequentially and their processing order is indicated by the red arrows.
II Related Work
Figure 2: Overview of the wavelet video coder pMCTF. (a): Encoding and decoding for a single temporal level with two frames. (b): Implementation details of the predict and update filters. $\bm{x}_{t}$ denotes the input video. $\bm{h}_{t}$ and $\bm{l}_{t}$ are the temporal highpass and lowpass subband.
II-A Learned Wavelet Image Coding
Ma et al. [10] introduced a learned wavelet image coding framework called iWave++ that follows the structure visualized in Fig. 1. The encoding and decoding transforms are implemented using the lifting scheme. They perform a forward and inverse wavelet transform based on Convolutional Neural Networks (CNNs). A context model considering dependencies between subbands and within the subband to be coded allows iWave++ to achieve state-of-the-art performance. There have been several extensions of iWave++ [11, 2, 3] including a parallelized context model in pWave++ [7]. Currently, there are ongoing efforts for standardizing wavelet image coders for the compression of regular images [12] and medical image 3D volumes [4].
II-B Learned Wavelet Video Coding
In the area of video coding, Dong et al. [13] proposed a partly trainable framework with a traditional temporal wavelet transform and a subband coder following iWave++. In [6], a fully trainable wavelet video coder based on MCTF was introduced. Its parallelized version pMCTF [7] serves as the baseline model for our investigations on end-to-end wavelet video coding.
II-B1 Model Overview - pMCTF
Fig. 2a provides an overview of pMCTF [7] for a single temporal decomposition level. A temporal wavelet transform realized via MCTF is followed by a spatial wavelet transform in pWave++, whereby all wavelet transforms are implemented using the lifting scheme. The lifting scheme consists of the three basic steps split, predict ($\mathcal{P}$) and update ($\mathcal{U}$). First, the input video sequence $\bm{x}_{t}$ is split into even and odd indexed frames $\bm{x}_{2t}$ and $\bm{x}_{2t+1}$. The temporal wavelet transform then computes a highpass and lowpass subband. The trainable predict step $\mathcal{P}$ performs a motion-compensated prediction to create the highpass subband containing residual information. The update step effectively performs lowpass filtering along the motion trajectory by reusing the computed motion vectors for inverse motion compensation on the highpass subband. The obtained temporal subbands are coded independently by dedicated pWave++ models with a parallelized four-step context model. The inverse temporal wavelet transform is obtained by reversing the lifting structure on the decoder side.
Figure 3: Coding a GOP of size 4 using MCTF. The temporal lowpass and highpass subbands are denoted as $\bm{l}_{t}^{(j)}$ and $\bm{h}_{t}^{(j)}$. Encoding starts at the top with the first decomposition level $j=1$. After every level finished frames can be transmitted, as indicated by gray coloring.
II-B2 Hierarchical GOP Structure in MCTF
MCTF relies on a dyadic decomposition that recursively applies a wavelet transform in temporal direction to the lowpass from the previous decomposition level. Hence, temporal scalability is provided with different temporal resolutions in different temporal layers. Fig. 3 exemplarily shows the structure of a GOP of size 4 with 2 temporal layers. In the first layer, the predict filter $\mathcal{P}^{(1)}$ is applied to the original frames $\bm{l}_{t}^{(0)}$. The residual highpass frames $h_{t}^{(1)}$ and associated motion vectors can be directly transmitted. The update step $\mathcal{U}^{(1)}$ gives the temporal lowpass frames $\bm{l}_{t}^{(1)}$ and the first level is finished. This temporal decomposition is repeated until there is only a single lowpass frame left. Every level has dedicated predict and update filters to account for different frame characteristics: $\mathcal{P}^{(1)}$ receives original frames an input, while one update step has been applied to the input of $\mathcal{P}^{(2)}$.
II-B3 Predict and Update Filters
Fig. 2b provides details on the implementation of the predict and update filters. For optical flow estimation, we use SpyNet [14] to obtain the motion vectors $\bm{m}_{t}$. For the motion vector compression module MV(j) we adopt the architecture from [15]. Based on the reconstructed motion vectors $\hat{\bm{m}}_{t}$, motion compensation (MC) is performed in the predict step and inverse motion compensation (MC-1) in the update step. Both the predict and update filter contain a trainable denoising module (DN), which follows a residual CNN structure. All trainable blue components in Fig. 2b depend on the temporal decomposition level $j$. Thus, there are dedicated modules for every temporal level.
However, pMCTF only supports training on two frames and different temporal levels are merely emulated by varying frame distances. This results in a train-test mismatch, because GOPs larger than two are used for testing and the hierarchical temporal layer structure can not be considered during training.
III Proposed Method
In the following, we introduce our proposed model which we refer to as pMCTF-L. Our model has three novel enhancements compared to the pMCTF model described above: Variable rate support, temporal layer adaptive quality scaling, and a multi-stage training strategy enabling training with multiple layers.
III-A Variable Rate Support
Up to now, learned MCTF coders [6, 7] require training dedicated models for every rate-distortion point. As popular approaches for continuous rate adaption like gain units [16] are not compatible with the wavelet latent space, we adopt a different approach based on [17]. That way, we are able to introduce a unified rate adaption mechanism to both motion vector coding and the temporal subband coder pWave++. We let our proposed model pMCTF-L support different Quantization Parameters (QPs) in a predefined range with $q_{\mathrm{num}}$ = 21 values in total. During training, we randomly choose a discrete quantization index $q$ for every batch element. During inference, the quantization index can be selected from the continuous range [0, 20]. We interpolate the rate-distortion trade off parameter $\lambda$ according to [17]:
$\lambda=\exp^{\ln\lambda_{\mathrm{min}}+\frac{q}{q_{\mathrm{num}}-1}\cdot(\ln\lambda_{\mathrm{max}}-\ln\lambda_{\mathrm{min}})},$ (1)
where $\lambda_{\mathrm{min}}$ and $\lambda_{\mathrm{max}}$ define a lower and upper bound for $\lambda$. In our implementation, we choose $\lambda_{\mathrm{min}}=0.03$ and $\lambda_{\mathrm{max}}=0.081$ to cover a wide quality range.
The quantization index $q$ determines the quantization process inside the learned wavelet video coder as well: The motion vector encoder and decoder have dedicated scaling parameters $s_{\mathrm{enc}}^{(j)}$ and $s_{\mathrm{dec}}^{(j)}$ for every temporal level (cf. Fig. 2b), while the pWave++ models have two quantization parameters QPL for the LL subband and QPH for the remaining subbands (cf. Fig. 2a). Both the layer-specific parameters for motion vector coding and the QPs of pWave++ are interpolated according to (1). Each parameter is within the range of an associated trainable lower and upper bound.
III-B Temporal Layer Adaptive Quality Scaling
Currently, learned MCTF coders do not consider their hierarchical temporal layer structure for bitrate allocation, even though this can give coding gains as known from traditional video coders. Therefore, we introduce an additional scaling parameter that depends on the decomposition level $j$ to allow adaptively scaling the quality depending on the temporal layer. We obtain a layer-dependent quantization parameter QPL for the lowpass subband and QPH for the highpass subbands as:
$\mathrm{QP}_{\mathrm{L}/\mathrm{H}}^{(j)}=\mathrm{QP}_{\mathrm{L}/\mathrm{H}}\cdot q_{\mathrm{scale}}^{(j)},$ (2)
with $j\in\left\{1,2,3\right\}$ for a GOP of size 8 with 3 temporal layers. Analogously to the quantization parameters, $q_{\mathrm{scale}}^{(j)}$ is in the range of a lower and upper bound defined by trainable parameters and interpolated based on the quantization index $q$ as specified in (1). In our experiments, we observe that $q_{\mathrm{scale}}^{(j)}$ increases the highpass quality with every decomposition level $j$. Hence, the highpass in the third decomposition level is coded with the highest quality. This behavior is sensible because this decomposition level influences the reconstruction of all frames from previous temporal levels.
III-C Multi-Stage Training Strategy
TABLE I: Proposed training schedule for a GOP size of 8. $d_{\mathrm{max}}$ denotes the maximum frame distance between frames in a training sample and $T$ the number of frames. ”All” trainable parts include the temporal wavelet transform with motion vector (MV) coding and the pWave++ models.
Stage Trainable parts $T$ $d_{\mathrm{max}}$ Loss Learning rate Epochs
[…]
Up to now, learned MCTF coders were only trained on two frames, corresponding to a single temporal layer. As training with multiple layers is not feasible from scratch, we propose a multi-stage training strategy that extends the schedules from [6] and [18]. Table I contains an overview of all six training stages. We initialize SpyNet with a pretrained model and keep the network fixed during all training stages, as we find this helps with overall training stability. Throughout all training stages, we randomly select a quantization index from the range $[0,q_{\mathrm{num}}-1]$ and use interpolated values for $\lambda$ as well as the quantization parameters described in Section III-A.
In the first two training stages (cf. Table I), only the motion vector coding modules are trainable. First, the loss is the distortion $D_{\mathrm{ME}}$ between the frame to be predicted $\bm{x}_{2t+1}$ and the prediction $\mathrm{MC}(\bm{x}_{2t})$ and then also contains the rate $R_{\mathrm{MV}}$ required for motion vector coding in the second stage. In the third training stage, all network components are trainable. The loss is the full rate-distortion loss for $T=2$ frames:
$\mathcal{L}_{\mathrm{full}}=\sum_{t=0}^{T-1}R_{\mathrm{all},t}+\lambda\cdot D_{\mathrm{MSE}}(\bm{x}_{t},\hat{\bm{x}}_{t}),$ (3)
where $t$ denotes the frame index and the distortion term corresponds to the Mean Squared Error (MSE) between the original frame $\bm{x}_{t}$ and the reconstructed frame $\hat{\bm{x}}_{t}$. $R_{\mathrm{all},t}$ consists of the rate required to code the temporal subbands using a pWave++ model and $R_{\mathrm{MV}}$ if the frame at index $t$ is a highpass.
In training stage 4 and 5, we select a random frame distance between one and $d_{\mathrm{max}}$ in every training step with equal probability. Depending on the frame distance, we now use different temporal wavelet transforms to model different temporal layers. For example, a frame distance of four corresponds to level $j=3$ with $\mathcal{P}_{3}$ and $\mathcal{U}_{3}$. The pWave++ models are kept fixed in stage 4 and are trainable again in stage 5. We tried omitting these training stages, however, we found that they are necessary as preparation for more temporal layers.
In the last stage 6, we use four frames per batch element to perform two full temporal decomposition levels. We randomly choose one of six combinations of frame indices with equal probability. With the first four frame index combinations $[i,i+1,i+2,i+3]$, where $i\in\left\{0,1,2,3\right\}$, we have a frame distance of one and use the temporal wavelet transform stages corresponding to the levels $j=1$ and $j=2$. For a frame distance of two, we have the index options $[0,1,3,5]$ or $[0,2,4,6]$ and we use the wavelet stages $j=2$ and $j=3$. Therefore, we effectively consider a GOP size of 8 with three levels. Additionally, we train the quality scaling parameter $q_{\mathrm{scale}}^{(j)}$ to allow quality adaption to different temporal layers.
IV Experiments and Results
IV-A Experimental Setup
IV-A1 Training
We use Vimeo90K [19] as training data set with a batch size of 8. During training, we crop patches of size $128\times 128$ from the luma channel of the respective training sample. We use AdamW as optimizer. We train pWave++ according to [7] for initializing the pWave++ models in our video coder. Different from [7], we incorporate variable rate support as described above.
IV-A2 Evaluation
We follow the test settings from [6, 20] and compare different coders on the first 96 frames of each test sequence. Our first test data set is the widely used UVG [21] data set containing seven sequences of size 1920$\times$1080. To consider a different resolution, we test on three sequences (CityAlley, FlowerFocus, FlowerKids) from the UVG 4K [21] data set with a resolution of 3840$\times$2160.
(a) UVG
(b) UVG 4K
(c) UVG - Ablation study
Figure 4: Rate-distortion evaluation and ablation study on different data sets.
IV-B Comparison to State-of-the-Art Video Coders
TABLE II: Rate-distortion evaluation on the UVG and UVG 4K data sets for a GOP size of 8. Average BD rate savings are provided relative to MCTF-CA [6] as anchor.
RGB-PSNR RGB-MS-SSIM
UVG UVG 4K UVG UVG 4K
DCVC-HEM [20] -8.44% -11.49% +6.66 % +22.88%
DCVC-DC [15] -15.06% -14.95% +1.63 % +21.00%
pMCTF-CA [7] +0.97% +81.45% +3.26% +91.38%
pMCTF-CA-L (ours) -17.49% -25.04% -8.73% -24.83%
VTM LD-P -91.36% -55.81% -10.68 % -31.89%
To evaluate state-of-the-art learned coders, we use DCVC-HEM [20] and DCVC-DC [15] with a GOP of 8 for a fair comparison to our MCTF approach. Because the GOP structure of MCTF is sensitive to the choice of GOP size, we allow a content adaptive GOP choice from [6] and obtain our model pMCTF-CA-L. In addition, we include other learned MCTF approaches in our experiments, namely MCTF-CA [6] and pMCTF-CA [7]. The latter has a parallelized context model in comparison to MCTF-CA, but both models have no variable rate support and were trained with a single temporal layer only. To consider a traditional hybrid video coder, we include VTM 17.2 [22]. We use VTM in Lowdelay P (LD-P) configuration with an intra period of 8 such that it performs unidirectional motion estimation and has the same I frame number as the learned coders.
The rate-distortion curves on the UVG data set can be seen in Fig. 4a. The curves demonstrate that our pMCTF-CA-L model performs especially well for high rates. Here, our approach is the only learning-based coder that outperforms VTM at bitrates above $\approx 0.11$ bpp. Because of the model’s capability of lossless compression, it has the capacity to perform specifically well for higher rates as well. In contrast, the models DCVC-HEM and DCVC-DC have better performance at low rates. For the UVG 4K data set, the rate-distortion curves in Fig. 4b show a similar behavior. Our pMCTF-CA-L model clearly outperforms the other MCTF models and also performs better than DCVC-HEM and DCVC-DC.
For a quantitative evaluation of the rate-distortion performance, we measure average Bjøntegaard Delta (BD) rate savings of different coders with MCTF-CA as an anchor. The results are shown in Table II. We find that our pMCTF-CA-L model outperforms the other learned coders in terms of PSNR as well as MS-SSIM on both the UVG and UVG 4K data set. In particular, we outperform DCVC-DC in terms of PSNR, which achieves BD rate savings of -15.06%, while our model gives savings of -17.49% on the UVG data set. On the UVG 4K data set, our model is about 10% better than DCVC-DC. The BD rate savings show that VTM outperforms the learning-based coders overall, whereby the performance gap is smaller for MS-SSIM.
IV-C Ablation Study
We demonstrate the effectiveness of our approach by an ablation study on the UVG data set. We take the pMCTF model from [7] as baseline. Fig. 4c provides the rate-distortion curves of the baseline pMCTF model and different configurations of our proposed method with BD rate savings in comparison to the baseline. The brown curve ”pMCTF var. rate” achieves average BD rate savings of -10.3% by enabling variable rate support, fixing the weights of the optical flow network, and using a four-step context model for motion vector coding. The model ”pMCTF var. rate + two lvls” gives BD rate savings of -16.5% by adding training stage six (cf. Table I) with two temporal levels. With the temporal layer adaptive quality scaling, we get the model ”pMCTF var. rate + two lvls + $q_{\mathrm{scale}}^{(j)}$” (pMCTF-L) achieving BD rate savings of -19.4% and hence further increasing coding efficiency. The red curve corresponding to pMCTF-CA-L with layer adaptivity and a content adaptive GOP choice achieves average BD rate savings of -23.0% over the baseline pMCTF model.
IV-D Complexity
TABLE III: Complexity comparison in terms of model size, kMAC/px, and decoding time. Runtime is measured on a NVIDIA RTX A6000 GPU on center crops of size $512\times 768$ from the UVG sequences.
Model Parameter size kMAC/px Decoding time
MCTF-CA [6] 90 MB 3554 1865.6s
pMCTF-CA [7] 182 MB 6593 7.7s
pMCTF-CA-L (ours) 144 MB 5357 8.7s
In Table III, we assess the complexity of our method compared to other MCTF-based approaches. To this end, we measure average decoding times, complexity in terms of kilo multiply-accumulate operations per pixel (kMAC/px), and provide the model size. Regarding the model size, the pMCTF models are larger due to the four-step context fusion models. We decreased the model size of our pMCTF-L compared to pMCTF by reducing the number of channels within the context fusion model. This results in less kMAC/px for our pMCTF-L model compared to pMCTF as well. In terms of kMAC/px, our model is more complex than the MCTF model from [6], however, the average decoding time is significantly lower, which makes our pMCTF-L model most practical. Also, our model only requires a single checkpoint for multiple rate-distortion points.
V Conclusion
In this paper, we introduced a training strategy that enables learned MCTF coders to consider two temporal layers during training. Thereby, we reduced the train-test gap and enabled temporal layer adpative quality scaling with flexible rate support. Overall, our proposed coder pMCTF-CA-L achieves BD rates savings of at least -17% compared to pMCTF-CA and also outperforms state-of-the-art learned coders like DCVC-DC. In future work, we will explore ways of considering information from multiple reference frames in the MCTF framework, e.g., by supporting bidirectional prediction.